OpenStudy (anonymous):

The centers of two 15.0 kg spheres are separated by 3.00 m. The magnitude of the gravitational force between the two spheres is approximately what? 1.11 × 10^-10 N 3.34 × 10^-10 N 1.67 × 10^-9 N 5.00 × 10^-9 N What is the answer and how did you get that answer? Please explain, thanks!

OpenStudy (s3a):

Main information: F_g = GmM/d^2 = (6.67×10^(-11))(15)(15)/(3)^2 F_g = 1.6675 × 10^(-9) F_g ≈ 1.67 × 10^(-9) N Background information: The variable d is the distance between the centres of mass which includes the radii of the two masses summed with the closest distance between their surfaces. Having said that, if the two masses are very far apart from each other as is the case with planets, the radii are negligible in lenth compared to the distance between the surfaces of the planets. So, in this problem, I assumed (and, I believe that the problem assumes) that the distance between the centres of mass of both spheres was 3m.

OpenStudy (anonymous):

Thank you so much! I got this question wrong before I posted it on here, but it makes sense now.

OpenStudy (s3a):

No problem. :)

OpenStudy (anonymous):

\[F=GM _{1}M _{2}divd ^{2}\]